OpenStudy (anonymous):

how do you find the bond order of the C-C and C-N bonds in the molecule CH3CN? (I know how to do this for diatomics but I don't know how in this scenario)

OpenStudy (chmvijay):

you want to find it based on molecular orbital diagram :) bond order is nothing but the no of bonds formed between those atoms

OpenStudy (anonymous):

so count the number of electrons in C-C bond (12) and use then fill the molecular orbital diagram and use the same technique as diatomics?

OpenStudy (chmvijay):

yup

OpenStudy (anonymous):

thanks! i'll give that a try!

OpenStudy (chmvijay):

Bond order =1/2 { no of electron in bonding orbitals- no of electron in anti bonding orbitals)

OpenStudy (anonymous):

\[(\sigma1s)^2(\sigma*1s)^2(\sigma2s)^2(\sigma*2s)^2(\pi*2p)4 \] so 8 in bonding orbitals - 4 in antibonding orbitals = 4 then divide by 2 = 2 the answer in the solution manual says it's 1. are they just simply asking for the the bond order between a C-C which is a single bond = 1?

OpenStudy (chmvijay):

yup may be

OpenStudy (anonymous):

okay got it, thank you! haha made an easy problem to complicated

OpenStudy (chmvijay):

i dont know exactly :) but it is not simple one as they have put how do you find ? in question

OpenStudy (kc_kennylau):